OpenStudy (anonymous):

HELP PLEASE sin[2arc cos (-3/5)] please show all the procedure

OpenStudy (alekos):

have they given you an answer?

OpenStudy (anonymous):

yes is \[-\frac{ 24 }{ 25 }\] but I don't know how to get there

OpenStudy (anonymous):

Okay. So when solving its quite simple... You just must know how to plug into the calculator So the answer is: \[-\frac{ 24 }{ 25 }\]

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @nicolepenafielm

OpenStudy (anonymous):

I need to solve this without the use of a calculator, thats why I am asking for help

OpenStudy (alekos):

let me help you because I just did it without a calculator

OpenStudy (anonymous):

thank you so much!

OpenStudy (alekos):

start with the expression inside the square brackets and let that equal \[\theta \] \[2arcos(-3/5) \]

OpenStudy (alekos):

so now we have \[2arcos(-3/5) = \theta \]

OpenStudy (alekos):

therefore \[arcos(-3/5) = \theta/2\]

OpenStudy (alekos):

so \[\cos(\theta/2) = -3/5\]

OpenStudy (alekos):

Now using a triangle we have the following...........